A 8.39 - nC charge is traveling at a constant velocity of 41900.00000000001 m/s near the surface of Earth, perpendicular to its magnetic field (B = 5 x 10-5 T). What is the mass of the charge?
Answer #1

as er equilibrium of forces

magnetic force = gravitational force

qVB = mg

8.39* 10^-9* 41900* 5* 10^-5 = m * 9.8

m = 1.909* 10^-9 kg

========

Comment before rate in case any doubt, will reply for sure.. goodluck